# Low-energy $N\phi$ scattering from a pole-enhanced triangle diagram

### v1

@@ Abstract
We investigate low-energy $N\phi$ scattering driven by a pole-enhanced triangle diagram, in which the two-kaon-exchange contribution is promoted by the near-threshold $\Lambda(1405)$ pole in the $N\bar{K}$ subsystem. Using an unphysical kaon mass motivated by lattice simulations, we evaluate the $N\phi$ scattering length and find that this mechanism generates an attractive interaction with a magnitude of 1.5-1.5 to $0.5,\rm{fm}$. Spin-dependent effects are not treated explicitly and are expected to provide subleading corrections in the near-threshold region. We further analyze the low-energy behavior of the triangle-diagram amplitude and show that the scattering length performs a characteristic power-law dependence on the parameter $\tilde{\delta}$, defined as the mass difference between the $K\bar{K}$ threshold and the $\phi$ meson. This threshold-driven behavior differs from that associated with van der Waals-type forces or the long-range tail of two-pion exchange, highlighting the role of three-body dynamics encoded in the pole-enhanced triangle diagram in shaping the near-threshold $N\phi$ interaction.
@@ I Introduction
At the loop level, contributions beyond tree-level meson exchange may also arise. By analogy with the two-pion-exchange mechanism studied in $J/\psi N$ scattering Fujii and Kharzeev (1999); Wu et al. (2024a), two-pion exchange has been considered in $N\phi$ scattering and incorporated into the interaction potential within the HAL QCD Lyu et al. (2022), following the structure discussed in Ref. Tarrús Castellà and Krein (2018). The two-pion-exchange contribution is evaluated through loop diagrams constructed from meson-baryon amplitudes. However, the leading-order Weinberg-Tomozawa interaction in chiral perturbation theory vanishes in $\phi\pi\to\phi\pi$ transition Roca et al. (2005). Therefor, two-pion exchange in $N\phi$ scattering originates from subleading chiral interactions and is expected to be suppressed in the low-energy region.
@@
Beyond the two-pion-exchange in $J/\psi N$ scattering, corrections from open-charmed hadron loops, represented by triangle diagrams, are considered in Ref. Brodsky and Miller (1997). Given that, the coupling of the $J/\psi$ to $D\bar{D}$ is suppressed compared to the $\phi K\bar{K}$ coupling, with a ratio $R_{1}\simeq 1/4$. In addition, the $DN$ interaction near threshold is weaker than the $\pi N$ interaction, leading to a ratio $R_{2}\simeq 0.1$. In the strange sector, these suppression factors are naturalness. The $\phi K\bar{K}$ coupling is unsuppressed, corresponding to $R_{1}\sim\mathcal{O}(1)$, while the Weinberg-Tomozawa interaction enhances the $N\bar{K}$ amplitude relative to $\pi N$ by a factor proportional to $m_{K}/m_{\pi}$. Furthermore, the proximity of the $\phi$ mass to the $K\bar{K}$ threshold induces a kinematical enhancement in the loop integral, which can be evaluated by a factor of $\left(m_{J/\psi}-2m_{D}\right)^{2}/\left(m_{\phi}-2m_{K}\right)^{2}$ when compared to the corresponding $D$-meson loop. These considerations indicate that triangle-diagram contributions associated with the $K\bar{K}$ threshold are enhanced in the $N\phi$ system. In particular, the proximity of the $K\bar{K}$ threshold, together with the presence of near-threshold structures in the $N\bar{K}$ subsystem, promotes the loop amplitude through both kinematics and dynamics. As a consequence, the triangle diagram provides an effective interaction with a range of $1/\sqrt{m_{K}(m_{\phi}-2m_{K})}$. Such loop-induced mechanisms are expected to play a non-negligible role in low-energy $N\phi$ scattering and contribute to the scattering length within the framework of the effective-range expansion. In addition, a characteristic threshold scaling behavior as a smoking-gun signal of three-body dynamics is encoded in the pole-enhanced triangle diagram.
@@
This article is organized as follows. In Sec. II, the effective-range expansion for low-energy $N\phi$ scattering associated with triangle-diagram contributions from two-kaon exchange is introduced. In Sec. III, the behavior of the $N\phi$ scattering length is presented and compared with results from HAL QCD and femtoscopic analyses by the ALICE collaboration. A summary and outlook are given in Sec. IV.
@@
where $\vec{\sigma}$ and $\vec{S}_{1}$ denote the Pauli matrices and the spin-1 operator, respectively, and $\vec{k}$ is the three-momentum of the nucleon in the center-of-mass frame. The $c_{0}$ represents the spin-independent (or spin-averaged) interaction, while $d_{1}^{\frac{1}{2},\,\frac{3}{2}}$ parameterizes the spin-dependent contribution for total spin $1/2$ and $3/2$, respectively.
@@
$\frac{8\pi\,th}{T}=V^{-1}-G=-\frac{1}{a^{\frac{1}{2},\,\frac{3}{2}}}-ik,$ (3)
@@
where $th=m_{N}+m_{\phi}$, $V$ denotes the effective potential, and $G=-\Lambda_{0}^{\frac{1}{2},\,\frac{3}{2}}/\pi+ik$ is the two-point loop function regularized by the cutoff $\Lambda_{0}$ Dong et al. (2021c). The corresponding scattering lengths are given by $a^{\frac{1}{2},\,\frac{3}{2}}=-1/\left(1/c_{0}+\Lambda_{0}^{\frac{1}{2},\,\frac{3}{2}}/\pi\right)$. When $\Lambda_{0}^{\frac{1}{2},\,\frac{3}{2}}$ are truncated at the same scale, which is concerned here, $a^{\frac{1}{2},\,\frac{3}{2}}$ are degenerated to be $a$, where the contribution from the spin-dependent interaction to the scattering length is attributed to be subleading Dong et al. (2021a, b) and not considered in the present study. The scattering length $a$ is a real quantity in a single-channel description, while it generally becomes complex once coupled-channel effects from inelastic channels below the $N\phi$ threshold are taken into account.
The $N\phi$ interaction can be conveniently decomposed into contributions from different dynamics,
$V=V_{\rm oth.}(\Lambda_{1})+V_{\rm tri.}(\Lambda_{2}),$ (4)
where $V_{\rm tri.}$ denotes the contribution associated with the triangle diagram, while $V_{\rm oth.}$ collects short-range interactions. The two terms are characterized by different renormalization scales, $\Lambda_{1}$ and $\Lambda_{2}$, reflecting their distinct dynamical origins. Here, $\Lambda_{1}$ represents a relatively hard scale associated with short-distance physics, whereas $\Lambda_{2}$ corresponds to a softer scale relevant for loop contributions near the $K\bar{K}$ threshold.
At high energy scale, generically denoted by $\Lambda_{1}$, short-range interactions contained in $V_{\rm oth.}$ are expected to provide the dominant interaction. We do not intend to model the full $N\phi$ interaction in a large energy range, but rather to isolate and quantify the threshold-driven dynamics that becomes prominent in the low-energy region around $\Lambda_{2}$. In this window, the triangle mechanism represents the leading nonperturbative contribution, while $V_{\rm oth.}$ is much smaller than the $V_{\rm tri.}$ and enters as a subleading short-distance correction, which can be referred to the renormalization group evolution Pavón Valderrama and Phillips (2015); Valderrama (2016); Peng et al. (2022) stemming from the Langer correction Langer (1937).
The relative size of short-range contributions can be obtained from scale separation and EFT power counting. The triangle diagram is characterized by a long-range scale $r_{\rm tri}\sim 1/\sqrt{m_{K}|\delta|}$ from the near-threshold Kaon loop, whereas short-range interactions are governed by $r_{\rm oth}\sim 1/\Lambda_{\chi}$ with $\Lambda_{\chi}\sim 1.0\,\rm GeV$. Since the scattering length is dominated by distances $r\sim 1/k_{\rm eff}$ and is governed by the spatial integral of the interaction, one expects
$\frac{a_{\rm oth}}{a_{\rm tri}}\sim\left(\frac{r_{\rm oth}}{r_{\rm tri}}\right)^{1}{\rm or}\left(\frac{r_{\rm oth}}{r_{\rm tri}}\right)^{3}$
implying a strong suppression of short-range interactions in the near-threshold regime, where the powers 1 and 3 correspond to the behavior of the triangle diagram in low-energy scattering. This is consistent with EFT power counting, where the triangle mechanism is infrared enhanced by the near-threshold momentum scale, while short-range interactions enter through local operators suppressed by the hard scale $\Lambda_{\chi}$. Thus, although short-distance physics can renormalize the overall strength, the threshold behavior of the scattering length is governed by the pole-enhanced triangle mechanism.
Such a separation of scales provides a useful framework for organizing the low-energy interaction, in which contributions from $V_{\rm oth.}$ are treated perturbatively around the threshold-sensitive dynamics encoded in $V_{\rm tri.}$. This hierarchy applies only within a narrow momentum window characterized by a low scale $\Lambda_{2}$, which is set by the proximity to the $K\bar{K}$ threshold and by the near-threshold structure in the $N\bar{K}$ subsystem.
@@
Figure 1: Schematic illustration of the two-kaon-exchange contribution to $N\phi$ scattering through a triangle diagram. The dashed, solid, and bold lines represent the kaon, $\phi$ meson, and nucleon, respectively. The corresponding particle mass and momentum are labeled in the figure.
@@
The term $V_{\rm tri.}$ corresponds to the two-kaon-exchange contribution represented by the triangle diagram shown in Fig. 1. This diagram involves two types of interaction vertices: (i) the coupling of the $\phi$ meson to a $K\bar{K}$ pair, and (ii) elastic $NK(\bar{K})$ scattering. The former is described by a vector-pseudoscalar-pseudoscalar interaction, with the effective Lagrangian given by
@@
The low-energy $NK$ and $N\bar{K}$ interactions entering the triangle diagram are estimated in chiral perturbation theory. At leading order, the vertex of the pseudoscalar meson scattering off the octet baryon is described by the Weinberg-Tomozawa interaction,
$\mathcal{L}_{WT}=\Big\langle\bar{B}\,i\gamma^{\mu}\frac{1}{4F_{\pi}^{2}}\,[\,[\Phi,\partial_{\mu}\Phi],\,B\,]\Big\rangle.$ (6)
where $\Phi$ and $B$ denote the SU(3) matrices for the pseudoscalar mesons and baryons, respectively,
$\Phi$ $=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{2}}\pi^{0}+\frac{1}{\sqrt{6}}\eta&\pi^{+}&K^{+}\\ \pi^{-}&-\frac{1}{\sqrt{2}}\pi^{0}+\frac{1}{\sqrt{6}}\eta&K^{0}\\ K^{-}&\bar{K}^{0}&-\frac{2}{\sqrt{6}}\eta\end{array}\right),$ (10)
$B$ $=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{2}}\Sigma^{0}+\frac{1}{\sqrt{6}}\Lambda&\Sigma^{+}&p\\ \Sigma^{-}&-\frac{1}{\sqrt{2}}\Sigma^{0}+\frac{1}{\sqrt{6}}\Lambda&n\\ \Xi^{-}&\Xi^{0}&-\frac{2}{\sqrt{6}}\Lambda\end{array}\right).$ (14)
This interaction leads to a low-energy structure of the form $\bar{u}\,\gamma_{\mu}(k_{1}^{\mu}+k_{2}^{\mu})\,u$ for the elastic $NK(\bar{K})$ channels, where $u$ and $\bar{u}$ are Dirac spinors and $k_{1}$ and $k_{2}$ denote the four-momenta of the incoming and outgoing mesons.
@@
The numerator of the loop amplitude $I(l)$ associated with the triangle diagram in Fig. 1 is written as
$\mathrm{Num}$ $=$ $g_{\rm eff}\,\epsilon^{\mu}\left(2l-P_{1}\right)_{\mu}\left(2l-P_{3}\right)_{\nu}\epsilon^{\nu}\,u(P_{2})\,\bar{u}(P_{4})$ (15)
$=$ $g_{\rm eff}\,\vec{\epsilon}\!\cdot\!\vec{\epsilon}^{\,\prime}\,\left(2\vec{l}-\vec{P}_{1}\right)\!\cdot\!\left(2\vec{l}-\vec{P}_{3}\right)\,u(P_{2})\,\bar{u}(P_{4})$
$=$ $g_{\rm eff}\,\left[4\,\vec{l}^{\,2}+\vec{P}_{1}\!\cdot\!\vec{P}_{3}\right]\,u(P_{2})\,\bar{u}(P_{4}),$
where $g_{\rm eff}$ denotes the product of the coupling constants at the interaction vertices. In non-relativistic $N\phi$ scattering, the intermediate kaons propagate close to their mass shell, such that $l^{0}-m_{K}\sim\vec{l}^{\,2}/(2m_{K})\ll m_{K}$.
@@
The loop amplitude is decomposed into partial waves of the $N\phi$ system. Near threshold, the S-wave component gives the leading contribution to the scattering amplitude, while higher partial waves are suppressed by additional powers of the external momenta. In particular, the D-wave contribution enters at higher order in the threshold expansion and does not contribute to the scattering length.
@@
$I(l)$ $=$ $\int\frac{d^{4}l}{(2\pi)^{4}}\frac{-g_{\rm eff}\left[4\,\vec{l}^{\,2}+\vec{P}_{1}\cdot\vec{P}_{3}\right]\,u(P_{2})\,\bar{u}(P_{4})}{\left(l^{2}-m_{1}^{2}\right)\left[\left(P_{1}-l\right)^{2}-m_{2}^{2}\right]\left[\left(P_{3}-l\right)^{2}-m_{3}^{2}\right]}$ (16)
$=$ $\int\frac{d^{4}l}{(2\pi)^{4}}\,\frac{-g_{\rm eff}}{8\,\omega_{1}\,\omega_{2}\,\omega_{3}}\,\frac{4\,\vec{l}^{\,2}+\vec{P}_{1}\cdot\vec{P}_{3}}{l^{0}-\omega_{1}+i\epsilon}\,\frac{u(P_{2})\,\bar{u}(P_{4})}{P_{1}^{0}-l^{0}-\omega_{2}+i\epsilon}\,\frac{1}{P_{3}^{0}-l^{0}-\omega_{3}+i\epsilon},$
where $m_{1}=m_{2}=m_{3}=m_{K}$ and the energies of the intermediate kaons are approximated as
$\omega_{1}$ $=$ $m_{K}+\frac{\vec{l}^{\,2}}{2m_{K}},$
$\omega_{2}$ $=$ $m_{K}+\frac{(\vec{P}_{1}-\vec{l})^{2}}{2m_{K}},$
$\omega_{3}$ $=$ $m_{K}+\frac{(\vec{P}_{3}-\vec{l})^{2}}{2m_{K}},$ (17)
@@
Near the $N\phi$ threshold, the intermediate kaons propagate close to their mass shell, such that $\omega_{1}-m_{K}\to 0$. The integrand of the loop function develops an $0/0$ structure. To extract the finite contribution in this limit, L’Hôpital’s rule is applied, leading to
$\frac{\omega_{1}-m_{K}}{\left(P_{1}^{0}-\omega_{1}-\omega_{2}\right)\left(P_{3}^{0}-\omega_{1}-\omega_{3}\right)}$ $=\frac{\omega_{1}-m_{K}}{\left(P_{1}^{0}-\omega_{2}\right)\left(P_{3}^{0}-\omega_{3}\right)+\omega_{1}^{2}-\omega_{1}\left(P_{1}^{0}+P_{3}^{0}-\omega_{2}-\omega_{3}\right)}$ (18)
$\xrightarrow{\omega_{1}\to m_{K}}{}\frac{1}{P_{1}^{0}+P_{3}^{0}-2\omega_{1}-\omega_{2}-\omega_{3}}\simeq\frac{1}{P_{1}^{0}-P_{3}^{0}-4\omega_{1}}.$
@@
$2\pi i\,\mathrm{Res}\left[I(z_{A})\right]=\int\frac{d^{3}l}{(2\pi)^{3}}\,\frac{g_{\rm eff}}{8\,m_{1}m_{2}m_{3}}\,\frac{8m_{K}}{P_{1}^{0}+P_{3}^{0}-4\omega_{1}}$
$+\int\frac{d^{3}l}{(2\pi)^{3}}\,\frac{g_{\rm eff}}{8\,m_{1}m_{2}m_{3}}\,\frac{\vec{P}_{1}^{\,2}}{P_{1}^{0}-\omega_{1}-\omega_{2}}\,\frac{1}{P_{3}^{0}-\omega_{1}-\omega_{3}},$ (19)

### v2

@@ Abstract
We investigate low-energy $N\phi$ scattering driven by a pole-enhanced triangle-like diagram, in which the two-Kaon-exchange contribution is promoted by the near-threshold $\Lambda(1405)$ pole in the $N\bar{K}$ subsystem. Using an unphysical Kaon mass motivated by lattice simulations, we evaluate the $N\phi$ scattering length and find that this mechanism generates an attractive interaction with a magnitude of 1.1-1.1 to $0.5\,\rm{fm}$. Spin-dependent effects are not treated explicitly and are expected to provide subleading corrections in the near-threshold region. We further analyze the low-energy behavior of the triangle-like diagram amplitude and show that the scattering length depends on the parameter $\delta$, defined as the mass difference between the $K\bar{K}$ threshold and the $\phi$ meson, and the pole position of $\Lambda(1405)$, where the $\Lambda(1405)$ plays a crucial role to understand $N\phi$ interaction. Furthermore, by employing physical hadron masses, our calculated scattering length is found to be consistent with current experimental data, providing a unified description across both unphysical and physical mass regimes. This type of interaction differs from that associated with van der Waals-type forces or the long-range tail of two-pion exchange, highlighting the role of three-body dynamics encoded in the pole-enhanced triangle-like diagram in shaping the near-threshold $N\phi$ interaction.
@@ I Introduction
At the loop level, contributions beyond tree-level meson exchange may also appear. By analogy with the two-pion-exchange mechanism investigated in $J/\psi N$ scattering Fujii and Kharzeev (1999); Wu et al. (2024a), two-pion exchange has been examined in $N\phi$ scattering and incorporated into the interaction potential within the HAL QCD Lyu et al. (2022), following the framework outlined in Ref. Tarrús Castellà and Krein (2018). The two-pion-exchange contribution is evaluated through loop diagrams constructed from meson-baryon amplitudes. However, the leading-order Weinberg-Tomozawa interaction in chiral perturbation theory vanishes in $\phi\pi\to\phi\pi$ transition Roca et al. (2005). Therefore, two-pion exchange in $N\phi$ scattering originates from subleading chiral interactions and is expected to be suppressed in the low-energy region.
@@
Beyond the two-pion-exchange in $J/\psi N$ scattering, corrections from open-charmed hadron loops, represented by triangle diagrams, are considered in Ref. Brodsky and Miller (1997). Given that, the coupling of the $J/\psi$ to $D\bar{D}$ is suppressed compared to the $\phi K\bar{K}$ coupling, with a ratio $R_{1}\simeq 1/4$. In addition, the $DN$ interaction near threshold is weaker than the $\pi N$ interaction, leading to a ratio $R_{2}\simeq 0.1$. In the strange sector, these suppression factors are naturalness. The $\phi K\bar{K}$ coupling is unsuppressed, corresponding to $R_{1}\sim\mathcal{O}(1)$, while the Weinberg-Tomozawa interaction enhances the $N\bar{K}$ amplitude relative to $\pi N$ by a factor proportional to $m_{K}/m_{\pi}$. Furthermore, the proximity of the $\phi$ mass to the $K\bar{K}$ threshold induces a kinematical enhancement in the loop integral, which can be evaluated by a factor of $\left(m_{J/\psi}-2m_{D}\right)^{2}/\left(m_{\phi}-2m_{K}\right)^{2}$ when compared to the corresponding $D$-meson loop. These considerations indicate that triangle-diagram contributions associated with the $K\bar{K}$ threshold are enhanced in the $N\phi$ system. In particular, the proximity of the $K\bar{K}$ threshold, together with the presence of near-threshold structures in the $N\bar{K}$ subsystem, promotes the loop amplitude through both kinematics and dynamics. As a consequence, the triangle diagram provides an effective interaction determined by $\sqrt{m_{K}(m_{\phi}-2m_{K})}$ and $\Lambda(1405)$. Such loop-induced mechanisms are expected to play a non-negligible role in low-energy $N\phi$ scattering and contribute to the scattering length within the framework of the effective-range expansion. In addition, a characteristic threshold scaling behavior as a smoking-gun signal of three-body dynamics is encoded in the pole-enhanced triangle diagram.
@@
This article is organized as follows. In Sec. II, the estimation of the $N\phi$ scattering length associated with triangle diagram contributions from two-Kaon exchange is introduced. In Sec. III, the behavior of the $N\phi$ scattering length is presented and compared with results from HAL QCD and femtoscopic analyses by the ALICE collaboration. A summary and outlook are given in Sec. IV.
@@
which is an analogy to the heavy hadron interaction Peng et al. (2022), where $\vec{\sigma}$ and $\vec{S}_{1}$ denote the Pauli matrices and the spin-1 operator, respectively, and $\vec{k}$ is the three-momentum of the nucleon in the center-of-mass frame. The $c_{0}$ represents the spin-independent (or spin-averaged) interaction, while $d_{1}^{\frac{1}{2},\,\frac{3}{2}}$ parameterizes the spin-dependent contribution for total spin $1/2$ and $3/2$, respectively.
@@
$\frac{8\pi\,th}{T}=\frac{1}{V+VGT}=-\frac{1}{a^{\frac{1}{2},\,\frac{3}{2}}}-ik,$ (3)
@@
where $th=m_{N}+m_{\phi}$, $V$ denotes the effective potential, and $G=-\Lambda_{0}^{\frac{1}{2},\,\frac{3}{2}}/\pi+ik$ is the two-point loop function regularized by the cutoff $\Lambda_{0}$ Dong et al. (2021c). When $\Lambda_{0}^{\frac{1}{2},\,\frac{3}{2}}$ are truncated at the same scale, which is concerned here, $a^{\frac{1}{2},\,\frac{3}{2}}$ are degenerated to be $a$, where the contribution from the spin-dependent interaction to the scattering length is attributed to be subleading Dong et al. (2021a, b) and not considered in the present study. To account for the off-shell configurations iterated through the two-point loop function $G$, the leading rescattering correction is evaluated. For a typical cutoff $\Lambda_{0}\sim 200\text{--}500~\text{MeV}$, the scattering length is modified to
$a_{N\phi}=V(1+\mathcal{E}_{\text{off}}),$ (4)
where the theoretical uncertainty induced by the off-shell effects is bounded by $\mathcal{E}_{\text{off}}\equiv|VGV/V|\simeq 8\%\text{--}23\%$. The smallness of this ratio indicates a rapid convergence of the iteration, ensuring that the effective potential $V$ evaluated on-shell robustly captures the dominant threshold dynamics.
@@
Figure 1: The two-Kaon-exchange contribution to $N\phi$ scattering through a triangle-like diagram. The dashed, solid, and thick lines represent the Kaon, $\phi$ meson, and nucleon, respectively. The corresponding particle mass and momentum are labeled in the figure.
@@
The term $V$ corresponds to the two-Kaon-exchange contribution represented by the triangle-like diagram shown in Fig. 1. This diagram involves two types of interaction vertices: (i) the coupling of the $\phi$ meson to a $K\bar{K}$ pair, and (ii) elastic $NK(\bar{K})$ scattering. The former is described by a vector-pseudoscalar-pseudoscalar interaction, with the effective Lagrangian given by
@@
The low-energy $NK$ and $N\bar{K}$ interactions entering the triangle diagram are estimated in chiral perturbation theory, where poles are dynamically generated in $N\bar{K}$ scattering. In the isoscalar channel, the $\Lambda(1405)$ resonance is widely interpreted as a dynamically generated $N\bar{K}$ bound state  Lu et al. (2023); Bulava and others (2024b, a); Guo et al. (2023); Zhuang et al. (2024); He et al. (2024). The proximity of this pole to the $N\bar{K}$ threshold leads to a significant enhancement of the low-energy $N\bar{K}$ amplitude, which in turn promotes the convergent part of the box-diagram contribution to $N\phi$ scattering. In the isovector $N\bar{K}$ channel, a virtual pole has been reported in several analyses and is sometimes associated with a broad $\Sigma$ excitation around $1.4~\mathrm{GeV}$ Ma and others (2023); Lu et al. (2023); Lyu et al. (2024); Wang et al. (2024); Li et al. (2024). This pole is located further from the physical sheet than the $\Lambda(1405)$ and therefore provides a subleading correction to the box diagram at threshold. By contrast, the $NK$ interaction in the $I=0$ and $I=1$ channels is governed by a vanishing ($C_{ij}=0$) and a repulsive ($C_{ij}=2$) Weinberg-Tomozawa term, respectively, and does not generate a near-threshold pole. As a result, the $NK$ subsystem does not induce an analogous enhancement of the convergent contribution. These considerations indicate that the near-threshold structure associated with the $\Lambda(1405)$ provides the leading enhancement mechanism in Fig. 1. The pole of $\Lambda(1405)$ reads
$t=-\lim_{s\rightarrow s_{R}}\frac{g_{i}^{2}}{s-s_{R}},$ (6)
with $N\bar{K}$ energy square $s$, pole position $\sqrt{s}_{R}=1417_{-4}^{+4}-i24_{-4}^{+7}\,\rm{MeV}$ and coupling $g_{i}=7.7_{-0.6}^{+1.2}\,\rm{GeV}$ Guo and Oller (2013). These parameters bridge the Fig. 1 and Fig. 2 (a).
To evaluate the $N\phi$ scattering length, the full calculation involves evaluating the diagrams in Fig. 2, which is very difficult, where the divergent parts in Fig. 2 (a) and (b) are regulator dependent. However, the energy corrections on the particles in the loop in Fig. 1 are small, and the pole of $\Lambda(1405)$ in Eq. (6) is approached by a propagator of $\Lambda(1405)$ in the box diagram in Fig. 3, where $s-s_{R}=2m_{\Lambda}\left(q^{0}_{\Lambda}-\omega_{\Lambda}\right)$.
Figure 3: Schematic illustration of the two-Kaon-exchange contribution to $N\phi$ scattering through a box diagram. The dashed, solid, thick, and thickened lines represent the Kaon, $\phi$ meson, nucleon, and $\Lambda(1405)$, respectively. The corresponding particle mass and momentum are labeled in the figure.
@@
Following the estimations of the vertices in the triangle-like diagram, the numerator of the loop amplitude $I(l)$ associated with the box diagram in Fig. 3 is written as
$\mathrm{Num}$ $=$ $-g_{\rm eff}\,\epsilon^{\mu}\left(2l-P_{1}\right)_{\mu}\left(2l-P_{3}\right)_{\nu}\epsilon^{\nu}\,u(P_{2})\,\bar{u}(P_{4})$ (7)
$=$ $-g_{\rm eff}\,\left[4\,\vec{l}^{\,2}+\vec{P}_{1}\!\cdot\!\vec{P}_{3}\right]\,u(P_{2})\,\bar{u}(P_{4}),$
@@
with polarization vector $\epsilon^{\mu}$, where $g_{\rm eff}$ denotes the product of the coupling constants at the interaction vertices. In non-relativistic $N\phi$ scattering, the intermediate Kaons propagate close to their mass shell, such that $l^{0}-m_{K}\sim\vec{l}^{\,2}/(2m_{K})\ll m_{K}$. The loop amplitude is decomposed into partial waves of the $N\phi$ system. Near threshold, the S-wave component gives the leading contribution to the scattering amplitude, while higher partial waves are suppressed by additional powers of the external momenta. In particular, the D-wave contribution enters at higher order in the threshold expansion and does not contribute to the scattering length.
@@
$I(l)$ $=$ $\int\frac{d^{4}l}{(2\pi)^{4}}\frac{-g_{\rm eff}\left[4\,\vec{l}^{\,2}+\vec{P}_{1}\cdot\vec{P}_{3}\right]\,u(P_{2})\,\bar{u}(P_{4})}{\left(l^{2}-m_{1}^{2}\right)\left[\left(P_{1}-l\right)^{2}-m_{2}^{2}\right]\left[\left(P_{3}-l\right)^{2}-m_{3}^{2}\right]\left[q^{2}_{\Lambda}-m_{\Lambda}^{2}\right]}$ (8)
$=$ $\int\frac{d^{4}l}{(2\pi)^{4}}\,\frac{-g_{\rm eff}}{8\,\omega_{1}\,\omega_{2}\,\omega_{3}}\,\frac{4\,\vec{l}^{\,2}+\vec{P}_{1}\cdot\vec{P}_{3}}{l^{0}-\omega_{1}+i\epsilon}\,\frac{u(P_{2})\,\bar{u}(P_{4})}{P_{1}^{0}-l^{0}-\omega_{2}+i\epsilon}\,\frac{1}{P_{3}^{0}-l^{0}-\omega_{3}+i\epsilon}\frac{1}{2m_{\Lambda}\left(q^{0}_{\Lambda}-\omega_{\Lambda}+i\,\epsilon\right)},$
where $m_{1}=m_{2}=m_{3}=m_{K}$ and the energies of the intermediate Kaons and $\Lambda(1405)$ are approximated as
$\omega_{1}$ $=$ $m_{K}+\frac{\vec{l}^{\,2}}{2m_{K}},\,\omega_{2}=m_{K}+\frac{(\vec{P}_{1}-\vec{l})^{2}}{2m_{K}},$
$\omega_{3}$ $=$ $m_{K}+\frac{(\vec{P}_{3}-\vec{l})^{2}}{2m_{K}},\,\omega_{\Lambda}=m_{\Lambda}+\frac{\left(\vec{P}_{1}+\vec{P}_{2}-\vec{l}\right)^{2}}{2m_{\Lambda}}$ (9)
@@
$2\pi i\,\mathrm{Res}\left[I(z_{A})\right]=\int\frac{d^{3}l}{(2\pi)^{3}}\,\frac{g_{\rm eff}}{8\,m_{K}^{3}}\,\frac{4\vec{l}^{2}}{P_{1}^{0}-\omega_{1}-\omega_{2}}\frac{u(P_{2})\,\bar{u}(P_{4})}{P_{3}^{0}-\omega_{1}-\omega_{3}}\frac{1}{2m_{\Lambda}\left(P_{1}^{0}+P_{2}^{0}-\omega_{1}-\omega_{\Lambda}\right)}$
$+\int\frac{d^{3}l}{(2\pi)^{3}}\,\frac{g_{\rm eff}}{8\,m_{K}^{3}}\,\frac{\vec{P}_{1}^{\,2}}{P_{1}^{0}-\omega_{1}-\omega_{2}}\,\frac{u(P_{2})\,\bar{u}(P_{4})}{2m_{\Lambda}\left(P_{1}^{0}+P_{2}^{0}-\omega_{1}-\omega_{\Lambda}\right)},$ (10)